write a mathematical equation for the following solutions and solve.
a number is 25 more than 2 times another number. if each of the numbers were multiplied by 5 the sum would be 350. what are the numbers? i have gotten to x is one number and y is another number so im at 2x+25=y and i dont understand what to do next

Answer by lwsshak3(11628)

write a mathematical equation for the following solutions and solve.
a number is 25 more than 2 times another number. if each of the numbers were multiplied by 5 the sum would be 350. what are the numbers? i have gotten to x is one number and y is another number so im at 2x+25=y and i dont understand what to do next
***
let x=1st number
2x+25=2nd number
5x+5(2x+25)=350 (if each of the numbers were multiplied by 5 the sum would be 350.
5x+10x+125=350
15x=225
x=15
2x+25=30+25=55
1st number=15
2nd number=55